Solve for s.
s^2 - 7s + 10 = 0

a. (2,5)
b. (-2,-5)
c. (-1,10)
d. (10,1)

Answer by checkley77(12844)

s^2 - 7s + 10 = 0
(s-5)(s-2)=0
s-5=0
s=5 ans.
s-2=0
s=2 ans.
Answer (2 & 5)

a. (2,5)
b. (-2,-5)
c. (-1,10)
d. (10,1)